40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5177     Entry time: Wed Aug 10 18:25:45 2011
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMY mini-update 

[Jenne, Jamie]

ETMY is now in its new nominal position, according to the rails that Kiwamu put in the other day.  OSEM voltages are all centered, and the magnets looked pretty well centered in the OSEM bores.  We're taking data for some free swinging spectra, to check the decoupling. 

Next up: Align Y-green to the arm, then move on to fixing the other optics that Jamie pointed out.

ELOG V3.1.3-